OpenStudy (anonymous):

to malevolence19: integral (dx)/[x(x^2 + 1)^2]

OpenStudy (anonymous):

Give me a sec to type it up :P

OpenStudy (jamesj):

You need to write out partial fractions: the integrand equals something/x + something/(x^2+1) + something/(x^2+1)^2 Let's see what mal19 comes up with ...

OpenStudy (anonymous):

i thought it would be partial fractions but the solutions manual has something completely different for the work, using x = tan(theta) dx = sec^2(theta)d(theta)

OpenStudy (anonymous):

Haha, well you can do it either way honestly. I would do it with partial fractions if it was me, but thats just me i suppose

OpenStudy (anonymous):

i'll try partial fractions, just the solutions manual threw me off when it used trig substitution

OpenStudy (jamesj):

oh, actually, the trig substitution here is easier

OpenStudy (anonymous):

You eventually have to use both. But I did partial fractions and then did trig sub. Tell me what you get and if you can't get it I'll help you. For partial fractions use this: \[\int\limits (original)=\int\limits \left[ \frac{A}{x}+\frac{Bx+C}{x^2+1}+\frac{Dx+E}{(x^2+1)^2} \right]dx\]

OpenStudy (jamesj):

dx/(x(x^2+1)^2) = du/(tan u.sec^2 u), where x = tan u and that equals = cos^3 u/sin u

OpenStudy (jamesj):

and that's straight forward, as cos^3 = cos - cos.sin^2

OpenStudy (anonymous):

Okay, after the trig sub (ignoring partial fractions) you should have (as james said)\[\int\limits \frac{\cos^3(\phi) d \phi}{\sin(\phi)}=\int\limits \frac{\cos(\phi)\cos^2(\phi)d \phi}{\sin(\phi)}=\int\limits \frac{\cos(\phi)(1-\sin^2(\phi))d \phi}{\sin(\phi)}\] Let u=sin(phi), du=cos(phi)dphi Then you have: \[\int\limits \frac{(1-u^2)du}{u}=\int\limits \frac{du}{u}-\int\limits u du=\ln|u|-\frac{1}{2}u^2+C\] But u=sin(phi) \[\ln|\sin(\phi)|-\frac{1}{2}\sin^2(\phi)+C\] Also: \[\sin(\phi)=\frac{x}{\sqrt{x^2+1}}\] So: \[\ln|\frac{x}{\sqrt{x^2+1}}|-\frac{1}{2} \left( \frac{x^2}{x^2+1} \right)+C\]